OpenStudy (anonymous):

1750 shared between 3 people. Ratio 3:5:8. What is the amount of the smallest portion. A population of 120 students is in the ratio of 1:9/ How many students are there to 4 teachers?

OpenStudy (campbell_st):

add the parts in the ratio... 3 + 5 + 8 = ...?

OpenStudy (anonymous):

ok. I did that. Then 1750/16. Got 109.37 Then multiplied by 3 Solution: 328.13.

OpenStudy (anonymous):

3x+5x+8x=1750

OpenStudy (anonymous):

Ok pgpilot326, I did that.

OpenStudy (campbell_st):

no just use 3 + 5 + 8 = 1 then to find the value of 1 unit in the ratio its 1750/16 = then multiply the answer by 3 to find the value of 3 part.

OpenStudy (campbell_st):

same for the other question... 1 + 9 = 10 120/10 = to find 1 prt... then multiply the answer by 4 to find the number of pupils for 4 teahcers

OpenStudy (anonymous):

the answer is 3x once you find x. x is the same thing as 1750/16... method matters not... go for understanding.

OpenStudy (anonymous):

Did you get 48 Campbell_st?

OpenStudy (anonymous):

@pgpilot326 : I got it thanks! My answer is 328.13.